a) Determine the market equilibrium price if the demand and supply function is given as:
D = 12p + 8
S = 14p – 4
Where D= demand
S=supply
p= price
b) Determine the equilibrium quantity if price is the same as above
D = 4p – 4q
S = 8q – 4p
Where D= demand
S=supply
p= price
q= quantity
a) Suppose the monthly income of an individual increases from Rs 20,000 to Rs 25,000 which increases his demand for clothes from 40 units to 60 units. Calculate the income elasticity of demand.
b) Quantity demanded for tea has increased from 300 to 400 units with an increase in the price of the coffee powder from Rs 25 to Rs 35. Calculate the cross elasticity of demand between tea and coffee.
1
Expert's answer
2015-08-30T12:10:17-0400
a) D = 12p + 8
S = 14p – 4
Equilibrium price is in the point, where Qd = Qs, so:
12p + 8 = 14p - 4
p = $6
b) The equilibrium quantity is in the point, where Qd = Qs, so if p = 6, then:
D = 4p – 4q
S = 8q – 4p
24 - 4q = 8q - 24
q = 4 units.
a) If income increases from Rs 20,000 to Rs 25,000 which increases his demand for clothes from 40 units to 60 units, then the income elasticity of demand is:
Ei = (60 - 40)/(25,000 - 20,000)*(25,000 + 20,000)/(60 + 40) = 20/5,000*45,000/100 = 1.8, so the good is normal.
b) If quantity demanded for tea has increased from 300 to 400 units with an increase in the price of the coffee powder from Rs 25 to Rs 35, then the cross elasticity of demand between tea and coffee is:
Ecp = (400 - 300)/(35 - 25)*(35 + 25)/(400 + 300) = 100/10*60/700 = 6/7, so the goods are substitutes.